OpenStudy (anonymous):

3|x − 8| + 15 = 18 What are the minimum and maximum temperatures for this day? x = 1, x = 17 x = −7, x = 9 x = 7 x = 9 No solution whats that awnser ? a,b,c,d?

OpenStudy (jdoe0001):

\(\bf 3|x - 8| + 15 = 18\implies 3|x - 8| =18-15\implies 3|x - 8| =3\\ \quad \\ |x - 8| =1\implies \begin{cases} +(x - 8) =1\\ \quad \\ \bf -(x - 8) =1 \end{cases}\)

OpenStudy (anonymous):

awnser this one too >>>>>Solve for x: −3|2x + 6| = −12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution a,b,c,d?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

63 and 12

OpenStudy (anonymous):

thats not a option...@Diamond-Co.OG.

OpenStudy (akashdeepdeb):

@seabass035 @jdoe0001 gave you the trick. Try solving that one yourself now. ^_^

OpenStudy (anonymous):

i dont understand anyof yhid so i cant @akashdeepdeb

OpenStudy (akashdeepdeb):

Whenever you see ANYTHING like this. |x| solutions can be +x or -x Because the very function/service of |...| is to give the ABSOLUTE or the POSITIVE Value. Getting this much?

OpenStudy (anonymous):

X=-1

OpenStudy (anonymous):

whiych one -1 or positive 1??@Madisaddledup

OpenStudy (anonymous):

negative because the end resault of the equation was -6x=6. when you divide a positive and a negative you get a what? negative.. so -1.

OpenStudy (anonymous):

thank you @Madisaddledup can u help memore?

OpenStudy (anonymous):

yeah, not right now though cuz i gotta do my own math:) haaha

OpenStudy (anonymous):

lol message me @Madisaddledup